OpenStudy (anonymous):

differentiate sin((1+x2)/(1-x2)), My question is that is a possible substitution possible ???

OpenStudy (amistre64):

differentiate means? find derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes...find d/dx of the function given...

OpenStudy (amistre64):

chain and quotient/product

OpenStudy (turingtest):

\[\frac{\sin(1+x^2)}{1-x^2}\]? and why do you want to substitute?

OpenStudy (amistre64):

i think the sin has the fraction as the argument :)

OpenStudy (anonymous):

its chain rule, \[\frac{dy}{dx} = \frac{dy}{du} \frac{du}{dx}\]

OpenStudy (anonymous):

\[\sin ((1+x2)\div(1-x2))\]

OpenStudy (anonymous):

so let u = (1+x^2)/(1-x^2) so y = sinu now find dy/du and du/dx and multiply to get dy/dx

OpenStudy (anonymous):

i know how to diferentiate it using quotient rule etc..but can i substitute any other function in place of x to make it simpler???

OpenStudy (anonymous):

@amistre64 - is it possible to substitute any other function in place of x to make it simpler???

OpenStudy (anonymous):

we dont tend to substiture for differentiation really, its integration where that comes in handy i mean substituting another function for x.. would it really be worth it? i cant see how it would make it simpler than just quotient ruling it, but if you have a substitution that does please share, i could be wrong

OpenStudy (amistre64):

x(x) is already a pretty simple function in its own right

OpenStudy (amistre64):

x=trig might produce something different tho ... not sure how simple it would make it tho

OpenStudy (anonymous):

suppose in this Q sin(inv)((1-x)/(1+x)) here we sub using cos2x... a similar sub i am asking for...

OpenStudy (amistre64):

x = cos2x is not a substitution .... x = cos(2t) might be

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

mistype